The simple interest on a certain sum of money for 3 years at 12% p.a is Rs 6750. What will be the compound interest on that sum at the same rate for the same period ?

Answer by satyareddy22(84)

S.I = (PTR/100)
S.I = 6750 ,T=3YRS,R=12%,p=?
SUBSTITUTING ALL VALUES IN THE ABOVE FORMULA
WE GET,P = 18750
NOW, A = P(1+R/100)^T
P=18750, T=3YRS,R = 12%
A= 18750(1+12/100)^3
= 18750(28/25)^3
=26342.4
NOW C.I = A - P
= 26342.4-18750=7592.4
SO,C.I = 7592.4